THE LOW-LYING CCN BENDING MODE OF THE CLOSE-TO-LINEAR MOLECULE NCCNO

\underline{HOLGER LICHAU}, MICHAEL LOCK, MANFRED WINNEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, D-35392 Gießen, Germany; STEPHEN C. ROSS, Department of Physics, University of New Brunswick, Fredericton NB E3B 5A3, Canada.

The band system of the low-lying CCN bending mode ν_{7} of the pseudohalofulminate NCCNO around 85 cm^{−1} has been recorded by high resolution FTIR spectroscopy to supplement our previous investigation of the pure rotational spectrum, which had revealed some slight, but significant indications of quasilinearity^{a}. Besides rovibrational transitions of the fundamental band, rovibrational transitions belonging to the first, second, third, and fourth hot band could be assigned. Using a conventional linear-molecule type Hamiltonian^{b}, these rovibrational transitions have been analyzed together with the pure rotational transitions. The results of this analysis confirm a moderate anharmonicity in the CCN bending potential. To quantify this anharmonicity, the combined data have been subjected to a semirigid bender analysis, which resulted in a rather flat bending potential with a significant quartic contribution.

aH. Lichau, B. P. Winnewisser, M. Winnewisser, C. W. Gillies, and J. Z. Gillies, 53^{rd} Int. Symposium, talk TH09 (1998)

bK. M. T. Yamada, F. W. Birss, and M. R. Aliev, J. Mol. Spectrosc. 112, 347–356 (1985)